geerky42 (geerky42):

\[\Large \int\limits_{1}^{2}(x-1)\sqrt{2-x}\space dx\]

OpenStudy (dumbcow):

im thinking integration by parts ... have you tried that?

geerky42 (geerky42):

wELL, i'M NEW TO INTEGRAL, i KNOW SOME TECHNIQUES BUT NOT COMPETELY CONFIDENT WITH IT. i'M CURRENTLY WORK WITH u-sUBSTITUTION. Sorry about caps.

geerky42 (geerky42):

I don't see how I can use u-substitution for this problem...

OpenStudy (dumbcow):

oh ok ... yeah substitution may work but i don't see it right away integration by parts is a technique when you have a product of 2 distinct functions \[\int\limits_{?}^{?} f(x) *g(x) dx\]

OpenStudy (dumbcow):

the equation is given as: \[\int\limits_{?}^{?} u*dv = uv -\int\limits_{?}^{?}v*du\] where u is f(x) and dv is g(x)

geerky42 (geerky42):

Ok, I'm giving it a try.

geerky42 (geerky42):

I don't think it will work. While I'm attempting to solve it, it's getting uglier.

OpenStudy (dumbcow):

no it will work but yes it can get uglier in the process ... anyway it takes some getting used to i found how u-substitution will work \[u = 2-x\] \[du = -dx\] \[\rightarrow -\int\limits_{?}^{?}(1-u) \sqrt{u} du = -\int\limits_{?}^{?} \sqrt{u} - u \sqrt{u}\]

geerky42 (geerky42):

Hmm clever! Thanks.

OpenStudy (dumbcow):

yw